# Exploratory Causal Inference in SAEnce

### v1

@@ 4 Neural Effect Search
To mitigate the multi-test significance collapse with entangled representation, we propose a novel causally principled algorithm that disentangles the leaked effects by recursive stratification:
@@
where NeuralEffectTest (Algorithm 2) is the procedure for multi-hypothesis testing on all the neurons $j$, by stratification (with arm-wise residualization) over the already retrieved effects S. See full description in Appendix B. The key idea is that if we test all neurons simultaneously, vanilla multiple testing cannot distinguish whether a neuron carries its own causal effect or merely leaks information about another concept. By contrast, NES first recovers the most prominent effect by its most representative neuron, then it controls its effect in subsequent tests, preventing the ECI paradox, continuing iteratively. Since this is a multiple testing setting, in Line 6 we still perform Bonferroni correction by dividing the significance level $\alpha$ by $m$. In practice, if the sample size is very small, one can make the test less conservative by relaxing the correction (which would return more, possibly false positives for the scientist to review).
@@
Discussion. NES recovers the $r$ effect concepts in probability and terminates, in sharp contrast with the paradox described earlier. While standard multi-hypothesis tests collapse with increasing power, i.e., $n$ and $\tau$, proposing all entangled neurons with $Y$ as significant effects, NES avoids this pitfall by recursively stratifying. Each iteration removes the spurious signal caused by leakage from already identified effects, bounding the collider bias, so that only the remain effect factor remains detectable. In this sense, NES does not merely test for effects: it disentangles the representation, identifying one true causal factor at a time until the entire effect subspace is recovered. Thus, NES can be interpreted both as a multiple-testing correction method robust to entanglement and as a principled effect disentanglement algorithm.
@@ 5 Related Works
Interpretable Heterogeneous Treatment Effect Estimation. A closely related line of work is the empirical discovery of treatment effect heterogeneity across covariates $W$. Methods such as causal trees, forests, and decision rules ensembles (Athey and Imbens, 2016; Athey et al., 2019; Bargagli-Stoffi et al., 2020) identify subpopulations with different responses, recognizing that pointwise estimation of the Conditional Average Treatment Effect (CATE) is almost impossible to test, and still difficult and risky to interpret. Since $W$ is lower-dimensional, interpretability of these partitions or rules is crucial, and the field has developed around making this empirical exploration scientifically meaningful. Our work is analogous in spirit: instead of asking who is affected (heterogeneity over $W$), we ask what is affected (discovering affected outcomes $Y$) when the outcome space itself is high-dimensional and initially unknown.
@@
Causal abstractions and representations. In the line of work of causal abstractions, Visual Causal Feature Learning (VCFL, Chalupka et al., 2014) was introduced to discover interventions in data rather than outcomes. In scientific trials, however, treatments are fixed by design, and the challenge is to recover their effects from complex outcome measurements. Causal Feature Learning (CFL, Chalupka et al., 2017) extends this to outcome clustering by grouping micro- to macro-variables using equivalence classes of $P(X\mid do(T))$. This requires density estimation in high-dimensional spaces, which is generally infeasible. While clustering other metrics may be possible, causal feature learning commits to a single grouping rule, while we find all statistically significant ones. Another line of work tackles the discovery of causal variables from high-dimensional observations (Schölkopf et al., 2021). Closest in spirit to our setting are interventional approaches (Varici et al., 2023, 2024; Zhang et al., 2023; Yao et al., 2025), which, even with all the necessary extra assumptions, would only offer identification results for the experimental settings $W$ and not the outcome (i.e., the component invariant to the intervention (Yao et al., 2025)). Therefore, they can not be applied to exploratory causal inference because they cannot discover outcome variables.
@@
Scientific discovery via SAEs. A related line of work uses SAEs to decompose polysemantic hidden representations in foundation models into more monosemantic units that align with single concepts (Bricken et al., 2023; Templeton et al., 2024; Huben et al., 2024; Papadimitriou et al., 2025). Although SAEs were initially proposed as an interpretability tool (Bricken et al., 2023), a growing body of negative results, including spurious interpretability on random networks (Heap et al., 2025), failures to isolated atomic concepts (Leask et al., 2025; Chanin et al., 2025), and limited downstream benefits (Wu et al., 2025), casts doubt on whether SAE features faithfully reflect underlying mechanisms rather than post-hoc artifacts. Despite these interpretability concerns, recent work shows that SAEs can still be useful for generating scientific hypotheses from high-dimensional data (Peng et al., 2025). For example, HypotheSAEs (Movva et al., 2025) leverage SAEs to surface human-understandable patterns correlated with target outcomes (e.g., engagement levels), which researchers can then treat as hypotheses for follow-up study. Our setting is related but distinct: whereas these approaches focus on correlations and do not provide statistical procedures to test the significance of the unsupervised discoveries, we target causal effects and develop inference to assess which high-dimensional outcomes $Y$ are affected, offering principled support for exploratory causal claims.
@@ 6.1 Semi-Synthetic Benchmark
We simulated a family of RCTs $\{T_{i},W_{i},Y_{i}\}_{i=1}^{n}$, relating both the individual covariates and outcomes one-to-one with specific attributes in the CelebA (Liu et al., 2018) dataset, e.g., wearing_hat and eyeglasses, and then assigned a random image $X_{i}$ from the dataset perfectly matching such attributes. Given the corresponding random sample $\{T_{i},X_{i}\}_{i=1}^{n}$ we (i) trained a SAE over the image representations encoded by SigLIP (Zhai et al., 2023), and (ii) tested NES for effect discovery against vanilla statistical tests ($t$-test, FDR (Schweder and Spjøtvoll, 1982), Bonferroni) and top-k effects selection. For quantitative evaluation, we first assessed SAE monosemanticity with respect to the considered CelebA attributes (see Figure 7), and extracted the ground truth neurons referring to $Y$. Then, for each effect discovery, we computed Recall, Precision, and Intersection over Union (IoU) with respect to them. Full details about the data generating procedure, training, and evaluation with additional assessment on interpretability and SAE entanglement, together with extensive ablations on method variants, i.e., estimator and test, and hedge cases, i.e., no-effect, are reported in Appendix D-E. The main results ($r=2$) are summarized in Figure 4.
@@ 6.2 Real-World Randomized Trial from Experimental Ecology
ISTAnt (Cadei et al., 2024) is an ecological experiment where ants from the same colony are randomly exposed to a treatment or a control substance and continuously filmed in triplets in a closed environment to study the concept of Social Immunity. The biologists are interested in identifying which latent behaviors are significantly affected by treatment. According to previous analysis, we first encoded each frame in the trial with DINOv2 (Oquab et al., 2023), and then we trained a SAE directly on the trial data. NES is then applied without Bonferroni adjustment due to the small sample size ($n=44$ videos) to discover treatment-sensitive codes, and only two neurons are returned.
Figure 5: Exploratory Causal Inference for Experimental Ecology. Without any knowledge of the behaviors of interest, our procedure retrieves two significant treatment effects, i.e., grooming and background, in agreement with previous literature.
Results. Figure 5 qualitatively summarizes the interpretations of such neurons, visualizing their corresponding most and least activated clips in the dataset. In agreement with the previous analysis on the dataset (Cadei et al., 2024, 2025), the first neuron retrieved (code 394) represents the grooming event, already measured as significantly affected by the treatment in any previous rationalist approach to the experiment, i.e., actually manually annotating and testing for it. Quantitatively, such a neuron is exactly the most predictive neuron for grooming event (F1-score=0.398) out of all the 4608 SAE’s codes, confirming the consistent results of our pipeline. We remark that our focus is on the identification of the effect as statistically significant. The imperfect F1-scorescore means that one should not compute treatment effects directly on the neural activation, e.g., without further labeling. The second neuron activated (code 550) represents the palette background (top right black color mark in the top left position in the first 4 batches of videos), which strongly correlates with the treatment due to the small experiment size (as discussed in the annotation bias by Cadei et al. (2024)). The fact that the model also identifies the effect of the treatment on the background due to the small sample size is a strength of the method: it is a statistically significant signal, and we want to retrieve it in addition to the behaviors since it is present in the dataset. Domain experts can select which signal is scientifically relevant and even use this information to improve their experimental settings.
@@ 7 Conclusion
In this paper, we have discussed how foundation models and SAEs can address the challenges of exploratory causal inference, serving as learned measurement devices. A key challenge is that SAE neurons may not map one-to-one onto high-level concepts, and even weak or mixed associations propagate the dependency on the treatment. This means that many neurons can be activated, making the interpretation difficult as they do not encode a single concept, and they activate more with larger sample sizes or stronger effects. We address this issue with Neural Effect Search, a statistical hypothesis testing procedure that iteratively controls for the biased dependency between neurons after they have been selected. Our experiments on semi-synthetic and real-world randomized trials are encouraging: our method uncovers both scientifically relevant effects and, when present, interpretable associations like background effects due to finite samples that experts can readily dismiss. Overall, we view this work as a first step toward AI-driven efficiency gains in exploratory data science, where foundation models can “look at massive amounts of data first” and then domain experts can identify which patterns have scientific value.
Our approach has several limitations. First, we strongly assume that the observed variables $X$ adequately capture information about the unknown $Y$, i.e., data sufficiency. For example, shrinkage in a tumor is detectable via X-ray imaging, but, depending on the tumor type, a treatment may also reduce its metabolic activity, which is measured with PET scans. More complex measurement processes for $X$, e.g., multi-modal are a natural extension of our work. Furthermore, we assume foundation models encode concepts linearly and that SAEs can approximately recover the effects. We believe the linear representation hypothesis (Park et al., 2023) is mild: even if current foundation models are imperfect, future iterations are likely to improve. The good “identifiability” assumption is our strongest. Promising early work already exists (Cui et al., 2025; Hindupur et al., 2025), but the identifiability theory of SAEs is not currently as well understood as that of causal representations (Yao et al., 2025), e.g., still unclear how to deal with continuous concepts. In our paper, we took a more empirical and future-looking stance on improvements in SAEs, focusing on inevitable finite samples entanglement while leveraging pretrained foundation models. Lacking identifiability means that domain experts can today only use our method “as a rescue system for hypotheses they may have missed”, before properly annotating the data and following the rationalist approach. We hope that our work will serve as a practical motivation for future work on identifiability in foundation model representations and SAE.
@@ Acknowledgments
We thank the Causal Learning and Artificial Intelligence group at ISTA for the continuous feedback on the project and valuable discussions. We further acknowledge the Fourth Bellairs Workshop on Causal Representation Learning held at the Bellairs Research Institute, February 14//21, 2025, and particularly the discussions with Chandler Squires, Dhanya Sridhar, Jason Hartford, and Frederick Eberhardt. Riccardo Cadei is supported by a Google Research Scholar Award and a Google Google-initiated gift to Francesco Locatello. Tommaso Mencattini was supported by the ISTernship program. This research was funded in part by the Austrian Science Fund (FWF) 10.55776/COE12). It was further partially supported by the ISTA Interdisciplinary Project Committee for the collaborative project “ALED” between Francesco Locatello and Sylvia Cremer. For open access purposes, the author has applied a CC BY public copyright license to any author-accepted manuscript version arising from this submission.
@@
All datasets used in this work are publicly available. In particular, the ISTAnt dataset (Cadei et al., 2024) was annotated and pre-processed by domain experts. While our model is capable of detecting statistically significant signals in randomized trials, the conclusions should not be interpreted as scientifically relevant unless domain experts interpret them. Since we cannot guarantee identifiability, it should only be used as a rescue system for hypotheses that may have been missed before committing to the rationalist approach, which is still necessary.
@@
Together with the paper, we submitted the code for NES, which can be used on top of any library for SAEs. A snipped Python implementation for the main algorithm is also presented in Appendix C. Standalone code to reproduce all experiments will be released with the final version of the paper. All the datasets we use are publicly available and experiment details are thoroughly detailed in Appendix D.
@@ Appendix A Proofs
Discussion. By the neural treatment effect decomposition, i.e., Equation 24, every code-level contrast lies in $\mathrm{span}\{v_{1},\dots,v_{r}\}$. Assumption A.1 guarantees this span is truly $r$-dimensional (nondegenerate) and that the linear map $\tau^{Y}\mapsto\tau^{Z}$ is injective—distinct effect vectors $\tau^{Y}$ produce distinct code-level contrasts. Informally, at the mean-effect level this behaves like “no loss of information” about $\tau^{Y}$ when passing through $V$ (akin to a sufficient statistic for $\tau^{Y}$ in a parametric family).
@@
Discussion. Assumption A.3 is a mean-level conditional requirement: conditioning within arms on the already-selected codes $Z_{\texttt{S}}$ explains their contribution in expectation, leaving only undiscovered effects in the adjusted contrast. In addition, Equation 30 is a mild pooled-strata transport bound: even though $G$ is post-treatment and may induce a collider opening, the weighted difference between observed and interventional arm means within strata is uniformly bounded by $\varepsilon$ and vanishing. In the linear (or locally linear) regime, the Jacobian of the mapping $Y\mapsto Z$ satisfies $\partial Z/\partial Y=V$ (or more generally $\partial\,\mathbb{E}[Z\mid Y,do(T)]/\partial Y=V$), so the columns $v_{k}$ are Jacobian columns; Assumption A.2 requires principal dominance, while Assumption A.3 ensures additive mean structure and bounded transport error so that stratification cancels already-discovered parts up to a uniform $\varepsilon$ bias.
@@
    Hence the corresponding $t$-statistics are $O_{p}(1)$ and, under Bonferroni control, no hypothesis is rejected with probability $\to 1$. Thus no further selections occur, and $\Pr\big(\texttt{S}_{\mathrm{final}}=\{j_{1},\dots,j_{r}\}\big)\to 1$ and $\mathbb{E}[|\texttt{S}_{\mathrm{final}}|]\to r.$
@@
Thus, for any stratification weights, the usual Neyman variance for the difference in means built on $R_{j,t}$ is weakly smaller asymptotically than that built on $Z_{j}$ (componentwise, within strata). Intuitively, residualization orthogonalizes $Z_{j}$ against already-discovered codes within arm, removing predictable variation and shrinking the standard error.
@@
and the Step 2 bounded-bias/cancellation proof applies with $R_{j}$ in place of $Z_{j}$. Hence residualization is asymptotically valid and (typically) more efficient. In small samples without splitting, we still target the same estimand in expectation up to $o_{p}(1)$ under standard regularity, but power can be non-monotone due to estimation noise.

### v2

@@ 4 Neural Effect Search
To mitigate the multi-test significance collapse with entangled representation, i.e., Paradox of Exploratory Causal Inference, we propose NeuralEffectSearch (NES), a novel causally principled algorithm to disentangle the leaked effects by recursive stratification:
@@
where NeuralEffectTest (Algorithm 2) is the procedure for multi-hypothesis testing on all the neurons $j$, by stratification (and potentially arm-wise residualization) over the already retrieved effects S. See Figure 4 for a simplified illustration of NES procedure, Appendix B for NeuralEffectTest description and Appendix C for a minimal Python implementation snippet. The key idea is that if we test all neurons simultaneously, vanilla multiple testing cannot distinguish whether a neuron carries its own causal effect or merely leaks information about another concept. By contrast, NES first recovers the most prominent effect by its most representative neuron, then it controls its effect in subsequent tests, preventing the ECI paradox, continuing iteratively. Since this is a multiple testing setting, in Line 6 we still perform Bonferroni correction by dividing the significance level $\alpha$ by $m$, preventing I type error, i.e., false neural effects discoveries. In practice, if the sample size is very small, one can make the test less conservative by relaxing the correction (which would return more, possibly false positives for the scientist to review).
(a) Estimate independently all the neural effects ($\tau_{Z_{1}}$$\tau_{Z_{2}}$$\tau_{Z_{3}}$); select the strongest ($\tau_{Z_{1}}$).
(b) Re-estimate the remaining neural effects controlling for the previous discovery ($Z_{1}$); select the strongest ($\tau_{Z_{3}}$).
(c) Iterate until no significant neural effects are measured, and return all the selected neurons ( $\text{{S}}=\{Z_{1},Z_{3}\}$).
Figure 4: Illustrative example of the Neural Effect Search (NES) recursive procedure with two latent variables $Y_{1}$, $Y_{2}$ affected by the treatment and three measurement channels $Z_{1}$, $Z_{2}$, $Z_{3}$, e.g., SAE codes, with $Z_{1}$ principally aligned with $Y_{1}$ and $Z_{3}$ with $Y_{2}$ (see principal alignment formal definition in Assumption A.2). At each iteration stratifying over the already selected neurons, e.g., $Z_{1}$ after the first iteration, acts as a proxy controlling for the mediated effects by the corresponding (true) latent effects already identified, i.e., $Y_{1}$.
@@
Discussion. NES recovers the $r$ effect concepts in probability and terminates, in sharp contrast with the paradox described earlier. While standard multi-hypothesis tests collapse with increasing power, i.e., $n$ and $\tau$, proposing all entangled neurons with $Y$ as significant effects, NES avoids this pitfall by recursively stratifying. Each iteration removes the spurious signal caused by leakage from already identified effects, and bounding the collider bias, only the unidentified effect factor remains detectable. In this sense, NES does not merely test for effects: it disentangles the effects (under Assumptions A.1-A.3), identifying and adjusting the estimation of one true treatment effect factor at a time until the entire effect subspace is recovered. Thus, NES can be interpreted both as a multiple-testing correction method robust to entanglement and as a principled effect disentanglement algorithm.
@@ 5 Related Works
Interpretable Heterogeneous Treatment Effect Estimation. A closely related line of work is the empirical discovery of treatment effect heterogeneity across covariates $W$. Methods such as causal trees, forests, and decision rules ensembles (athey2016recursive; athey2019generalized; bargagli2020causal) identify subpopulations with different responses, recognizing that pointwise estimation of the Conditional Average Treatment Effect (CATE) is almost impossible to test, and still difficult and risky to interpret. Since $W$ is lower-dimensional, interpretability of these partitions or rules is crucial, and the field has developed around making this empirical exploration scientifically meaningful. Our work is analogous in spirit: instead of asking who is affected (heterogeneity over $W$), we ask what is affected (discovering affected outcomes $Y$) when the outcome space itself is high-dimensional and initially unknown.
@@
Causal abstractions and representations. In the line of work of causal abstractions, Visual Causal Feature Learning (VCFL, chalupka2014visual) was introduced to discover interventions in data rather than outcomes. In scientific trials, however, treatments are fixed by design, and the challenge is to recover their effects from complex outcome measurements. Causal Feature Learning (CFL, chalupka2017causal) extends this to outcome clustering by grouping micro- to macro-variables using equivalence classes of $P(X\mid do(T))$. This requires density estimation in high-dimensional spaces, which is generally infeasible. While clustering other metrics may be possible, CFL commits to a single grouping rule, while we find all statistically significant ones. Another line of work tackles the discovery of causal variables from high-dimensional observations (scholkopf2021toward). Closest in spirit to our setting are interventional approaches (varici2023score; varici2024general; zhang2023identifiability; yaounifying), which, even with all the necessary extra assumptions, would only offer identification results for the experimental settings $W$ and not the outcome (i.e., the component invariant to the intervention (yaounifying)). Therefore, they can not be applied to exploratory causal inference because they cannot discover outcome variables.
@@
Scientific discovery via SAEs. A related line of work uses SAEs to decompose polysemantic hidden representations in foundation models into more monosemantic units that align with single concepts (bricken2023monosemanticity; templeton2024scaling; huben2024sparse; papadimitriou2025interpretinglinearstructurevisionlanguage). Although SAEs were initially proposed as an interpretability tool (bricken2023monosemanticity), a growing body of negative results, including spurious interpretability on random networks (heap2025sparseautoencodersinterpretrandomly), failures to isolated atomic concepts (leask2025sparse; chanin2025a), and limited downstream benefits (wu2025axbench), casts doubt on whether SAE features faithfully reflect underlying mechanisms rather than post-hoc artifacts. Despite these interpretability concerns, recent work shows that SAEs can still be useful for generating scientific hypotheses from high-dimensional data (peng2025usesparseautoencodersdiscover). For example, HypotheSAEs (movva2025sparse) leverage SAEs to surface human-understandable patterns correlated with target outcomes (e.g., engagement levels), which researchers can then treat as hypotheses for follow-up study. Our setting is related but distinct: whereas these approaches focus on correlations and do not provide statistical procedures to test the significance of the unsupervised discoveries, we target causal effects and develop inference to assess which high-dimensional outcomes $Y$ are affected, offering principled support for exploratory causal claims.
@@ 6.1 Semi-Synthetic Benchmark
We simulated a family of RCTs $\{T_{i},W_{i},Y_{i}\}_{i=1}^{n}$, relating both the individual covariates and outcomes one-to-one with specific attributes in the CelebA (liu2018large) dataset, e.g., wearing_hat and eyeglasses, and then assigned a random image $X_{i}$ from the dataset perfectly matching such attributes. Given the corresponding random sample $\{T_{i},X_{i}\}_{i=1}^{n}$ we (i) trained a SAE over the image representations encoded by SigLIP (zhai2023sigmoid), and (ii) tested NES for effect discovery against vanilla statistical tests ($t$-test, FDR (schweder1982plots), Bonferroni) and top-$k$ effects selection. For quantitative evaluation, we first assessed SAE monosemanticity with respect to the considered CelebA attributes (see Figure 8), and extracted the ground truth neurons referring to $Y$. Then, for each effect discovery, we computed Recall, Precision, and Intersection over Union (IoU) with respect to them. Full details about the data generating procedure, training, and evaluation are reported in Appendix D, and the main results ($r=2$) are summarized in Figure 5. Additional experiments testing (i) the required assumptions, (ii) method consistency with extensive ablations, and (iii) additional baselines are reported in Appendix E.
@@ 6.2 Real-World Randomized Trial from Experimental Ecology
ISTAnt (cadei2024smoke) is an ecological experiment where ants from the same colony are randomly exposed to a treatment or a control substance and continuously filmed in triplets in a closed environment to study the concept of Social Immunity. The biologists are interested in identifying which latent behaviors are significantly affected by treatment. According to previous analysis, we first encoded each frame in the trial with DINOv2 (oquab2023dinov2), and then we trained a SAE directly on the trial data. NES is then applied without Bonferroni adjustment due to the small sample size ($n=44$ videos) to discover treatment-sensitive codes, and only two neurons are returned.
Figure 6: Exploratory Causal Inference for Experimental Ecology. Without any knowledge of the behaviors of interest, our procedure retrieves two significant treatment effects, interpreted as (i) grooming behavior and (ii) finite sample experimental design error, in full agreement with previous literature.
Results. Figure 6 shows a qualitative interpretation of such neurons, by filtering the most and least activated clips in the dataset. In agreement with the previous analysis on the dataset (cadei2024smoke; cadei2025causal), the first neuron retrieved (code 394) seems to separate grooming events, already measured as significantly affected by the treatment in any previous rationalist approach to the experiment, i.e., actually manually annotating and testing for it. Quantitatively, such a neuron is exactly the most predictive neuron for grooming event (F1-score=0.398=0.398) out of all the 4 608 SAE’s codes, confirming the consistent results of our pipeline. The imperfect F1-$score<1$ suggests the possibility of other entangled effects or broader representation, prohibiting linking the treatment effect on the interpreted outcome with the effect on such a neuron without further labeling. The second neuron activated (code 550) represents a black position marking in the background, i.e., a specific black color marking for dish palette positioning on the background (top right outline). Indeed, all the videos recorded with such visible position marking (top left dish palette in the first 4 batches of recordings) are always assigned to the same treatment by chance due to the small experiment size, as discussed in the annotation bias by cadei2024smoke. Quantitatively, it is exactly the most predictive neuron for such a characteristic (F1-$score=0.568$). The fact that the model also identifies the effect of the treatment assignment on the background due to the small sample size is a strength of the method: it is still a statistically significant signal and should not be ignored, even if not interesting directly as a biological effect. It is indeed the domain experts’ duty to select which signal is scientifically relevant or just useful to eventually (not necessarily) improve the experiment design.
@@ 7 Conclusion
In this paper, we have discussed how foundation models and SAEs can address the challenges of exploratory causal inference, serving as learned measurement devices. A key challenge is that SAE neurons may not map one-to-one onto high-level concepts, and even weak or mixed associations propagate the dependency on the treatment. This means that many neurons can be activated, making the interpretation difficult as they do not encode a single concept, and they activate more with larger sample sizes or stronger effects. We address this issue with Neural Effect Search, a statistical hypothesis testing procedure that iteratively controls for the biased dependency between neurons after they have been selected. Our experiments on semi-synthetic and real-world randomized trials are encouraging: our method uncovers both scientifically relevant effects and, when present, interpretable finite sample associations, e.g., background marking, that experts can readily dismiss. Overall, we view this work as a first step toward AI-driven efficiency gains in exploratory data science, where foundation models can “look at massive amounts of data first” and then domain experts can identify which patterns have scientific value.
Our approach has several limitations. First, we strongly assume that the observed variables $X$ adequately capture information about the unknown $Y$, i.e., data sufficiency. For example, shrinkage in a tumor is detectable via X-ray imaging, but, depending on the tumor type, a treatment may also reduce its metabolic activity, which is measured with PET scans. More complex measurement processes for $X$, e.g., multi-modal are a natural extension of our work. Furthermore, we assume foundation models encode concepts linearly and that SAEs can approximately recover the effects. We believe the linear representation hypothesis (park2023linear) is mild: even if current foundation models are imperfect, future iterations are likely to improve. The good “identifiability” assumption is our strongest. Promising early work already exists (cui2025theoretical; hindupur2025projecting), but the identifiability theory of SAEs is not currently as well understood as that of causal representations (yaounifying), e.g., still unclear how to deal with continuous concepts. In our paper, we took a more empirical and future-looking stance on improvements in SAEs, focusing on inevitable finite samples entanglement while leveraging pretrained foundation models. Lacking identifiability means that domain experts can today only use our method “as a rescue system for hypotheses they may have missed”, before properly annotating the data and following the rationalist approach. We hope that our work will serve as a practical motivation for future work on identifiability in foundation model representations and SAE.
@@ Acknowledgments
We thank the Causal Learning and Artificial Intelligence group at ISTA for the continuous feedback on the project and valuable discussions. We further acknowledge the Fourth Bellairs Workshop on Causal Representation Learning held at the Bellairs Research Institute, February 14//21, 2025, and particularly the discussions with Chandler Squires, Dhanya Sridhar, Jason Hartford, and Frederick Eberhardt and further feedback from Judea Pearl. Riccardo Cadei is supported by a Google Research Scholar Award and a Google Google-initiated gift to Francesco Locatello. Tommaso Mencattini was supported by the ISTernship program. This research was funded in part by the Austrian Science Fund (FWF) 10.55776/COE12). It was further partially supported by the ISTA Interdisciplinary Project Committee for the collaborative project “ALED” between Francesco Locatello and Sylvia Cremer. For open access purposes, the author has applied a CC BY public copyright license to any author-accepted manuscript version arising from this submission.
@@
All datasets used in this work are publicly available. In particular, the ISTAnt dataset (cadei2024smoke) was annotated and pre-processed by domain experts. Despite our method guarantees identification under untestable representation learning assumption (without supervision), any conclusion should not be interpreted as scientifically relevant without a domain expert’s interpretation. Additionally, since we cannot guarantee identifiability, it should only be used as a rescue system for hypotheses that may have been missed before committing to the rationalist approach, which is anyway necessary for inference, i.e., treatment effect estimation and uncertainty quantification.
@@
All the datasets we use are publicly available and experiment details are thoroughly detailed in Appendix D-E. A snipped Python implementation for the main algorithm is also presented in Appendix C.
@@ Appendix A Proofs
Discussion. By the neural treatment effect decomposition, i.e., Equation 24, every code-level contrast lies in $\mathrm{span}\{v_{1},\dots,v_{r}\}$. Assumption A.1 guarantees this span is truly $r$-dimensional (nondegenerate) and that the linear map $\tau^{Y}\mapsto\tau^{Z}$ is injective—distinct effect vectors $\tau^{Y}$ produce distinct code-level contrasts. Informally, at the mean-effect level, this behaves like “no loss of information” about $\tau^{Y}$ when passing through $V$ (akin to a sufficient statistic for $\tau^{Y}$ in a parametric family).
@@
Discussion. Assumption A.3 is a mean-level conditional requirement: conditioning within arms on the already-selected codes $Z_{\texttt{S}}$ explains their contribution in expectation, leaving only undiscovered effects in the adjusted contrast. In addition, Equation 30 is a mild pooled-strata transport bound: even though $G$ is post-treatment and may induce a collider opening, the weighted difference between observed and interventional arm means within strata is uniformly bounded by $\varepsilon$ and vanishes. In the linear (or locally linear) regime, the Jacobian of the mapping $Y\mapsto Z$ satisfies $\partial Z/\partial Y=V$ (or more generally $\partial\,\mathbb{E}[Z\mid Y,do(T)]/\partial Y=V$), so the columns $v_{k}$ are Jacobian columns; Assumption A.2 requires principal dominance, while Assumption A.3 ensures additive mean structure and bounded transport error so that stratification cancels already-discovered parts up to a uniform $\varepsilon$ bias.
@@
Note. Since the Theorem establishes $\Pr\!\big(\texttt{S}_{\mathrm{final}}=\{j_{1},\dots,j_{r}\}\big)\to 1$, it follows directly that also the overall family-wise error rate (FWER) and false discovery rate (FDR) of the entire recursive NES procedure converge to 0.
@@
    Hence, the corresponding $t$-statistics are $O_{p}(1)$ and, under Bonferroni control, no hypothesis is rejected with probability $\to 1$. Thus, no further selections occur, and $\Pr\big(\texttt{S}_{\mathrm{final}}=\{j_{1},\dots,j_{r}\}\big)\to 1$ and $\mathbb{E}[|\texttt{S}_{\mathrm{final}}|]\to r.$
@@
Convergence Rate
The proof of Theorem 4.1 shows that, at each round and for each remaining principal neuron $j_{k}$, the corresponding $t$-statistic has a noncentrality parameter of order $\sqrt{n}$. In particular, the smallest principal margin that NES can detect with nontrivial power scales as $O(n^{-1/2})$, as in standard CLT-based tests.
@@
Thus, for any stratification weights, the usual Neyman variance for the difference in means built on $R_{j,t}$ is weakly smaller asymptotically than that built on $Z_{j}$ (componentwise, within strata). Intuitively, residualization orthogonalizes $Z_{j}$ against already-discovered codes within the arm, removing predictable variation and shrinking the standard error.
@@
and the Step 2 bounded-bias/cancellation proof applies with $R_{j}$ in place of $Z_{j}$. Hence, residualization is asymptotically valid and (typically) more efficient. In small samples without splitting, we still target the same estimand in expectation up to $o_{p}(1)$ under standard regularity, but power can be non-monotone due to estimation noise.